OpenStudy (anonymous):

How do you simplify this rational expression?:

OpenStudy (anonymous):

\[\frac{ 3x+2 }{ 15x ^{2}+16x+27 }\]

OpenStudy (anonymous):

you don't?

OpenStudy (anonymous):

It says you can?

OpenStudy (anonymous):

3x + 2 is not a factor of 15x^2 + 16x + 27

OpenStudy (anonymous):

What about \[\frac{ x ^{2}+6x+9 }{ x ^{2}+12x+27 }\]

OpenStudy (anonymous):

I typed it wrong(sorry)

OpenStudy (anonymous):

\[\frac{x^2 + 6x + 9}{x^2 + 12x + 27} = \frac{(x + 3)(x + 3)}{(x + 3)(x + 9)}\]

OpenStudy (anonymous):

you can factor out the (x + 3)s

OpenStudy (anonymous):

and say that it is (x+3)/(x+9)

OpenStudy (anonymous):

I guess not so much factor out as really just cancel

OpenStudy (anonymous):

and I was talking about one (x+3) on the top and one (x+3) on the bottom

OpenStudy (anonymous):

So the expression would be: \[\frac{ (x+3) }{ (x+9) }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

you are welcome